A sociology professor assigns letter grades on a test according to the following scheme. A: Top 6% of scores B: Scores below the top 6% and above the bottom 64% C: Scores below the top 36% and above the bottom 15% D: Scores below the top 85% and above the bottom 6% F: Bottom 6% of scores Scores on the test are normally distributed with a mean of 67.7 and a standard deviation of 9.3. Find the minimum score required for an A grade. Round your answer to the nearest whole number, if necessary.
Answer #1

Ans:

mean =67.7

standard deviation = 9.3

For A: Top 6% of scores

P(Z>z)=0.06

P(Z<=z)=1-0.06=0.94

z=normsinv(0.94)

z=1.555

Minimum score for an A grade=67.7+1.555*9.3=82.16